OpenStudy (anonymous):

How do you evaluate sin(arctan -4/3)?

OpenStudy (anonymous):

Find arctan of -4/3 first. Drawing the triangle on the unit circle is good for this.

OpenStudy (anonymous):

I know how to solve for arctan positive 4/3 but I'm not sure how to solve it when it is negative.

OpenStudy (anonymous):

tangent is opposite over adjacent, which means that to get -4/3, either the opposite side or the adjacent side must be negative. 

OpenStudy (anonymous):

I think because of domain and range issues, you need to take the fourth quadrant angle.

OpenStudy (calculusfunctions):

Right. Only the triangle in the fourth quadrant need be considered because the inverse tangent function is defined over the interval\[\frac{ - \pi }{ 2 }<x <\frac{ \pi }{ 2 }\]

OpenStudy (calculusfunctions):

 By Pythagorean theorem the third side (hypotenuse side) is 5. Now what is the sine ratio of the angle the terminal arm forms with the x-axis? That ratio is your desired answer.